OpenStudy (anonymous):

Use synthetic division to determine if the the given value for k is a zero of the corresponding polynomial. If not, determine p(k).

OpenStudy (anonymous):

\[P (x)= x^4 -1; k=\]

OpenStudy (anonymous):

K=1

OpenStudy (anonymous):

Can you help me on this which equation represents the line that passes thou the point (1,5) and had a slope of (-2)

OpenStudy (anonymous):

These are the choices 1- y= 2x+72- y= -2x+113- y=2x-94-y=2x+3

hero (hero):

@Kumar2006 , you should show proper manners by not posting your question in other people's questions.

OpenStudy (anonymous):

I think it is the second one

OpenStudy (anonymous):

Is it or no?

OpenStudy (campbell_st):

use the remainder theorem to check for a remainder... if there is no remainder then x-1 is a factor P(1) = 1^4 -1 = 0 then x - 1 is a factor the polynomial is P(x) = (x^2 -1)(x^2 + 1) = (x-1)(x+1)(x^2+ 1)

OpenStudy (anonymous):

What is the answer?